# Multiplicity of young isolated planetary mass objects in Taurus and Upper Scorpius

### v1

@@ 5.3 Implications for the formation process of FFPs
However, this simple picture–where high stellar density and feedback from massive stars prevent or disrupt binary FFPs–becomes more complex and puzzling when considering the discoveries of wide binary FFP candidates in several young, nearby massive clusters. Luhman (2024) reported the detection of one candidate binary FFP ([LAB2024] 139) among a sample of 242 BDs and FFPs identified in the 1 Myr, 400 pc Orion Nebula Cluster (ONC). This system exhibits a wide separation of 132 au and a relatively small mass ratio of q0.13. Similarly, Langeveld et al. (2024) report two binary candidates among their 6 FFP candidates in the 1–3 Myr, 300 pc NGC 1333 cluster (NGC1333-NN10 and NGC1333-NN12). Both systems have wide separations (164 au and 76 au, respectively) but relatively ”common” nearly equal mass ratios of 0.7–0.8.
If confirmed as physical pairs, these 3 wide binary candidates with primaries at or near the planetary mass regime would join the two previously known young wide binaries Oph 11 ( au, Close et al., 2007) and Oph 98, ( au, Fontanive et al., 2020) located in the Ophiuchus star forming region near the Oph pair of early-B stars, as well as the 2 000 au system LOri 167 made of a 18 MJup and 8MJup pair in the Collinder 69 cluster near the massive Ori pair of OB stars (Barrado Y Navascués et al., 2007), as illustrated in Fig. 11. They would also add to the similarly weakly bound wide binary BD candidates reported in Bouy et al. (2006) and Béjar et al. (2008) in USco. Notably, only one such wide binary BD has been reported to date in Taurus (FU Tau B, Luhman et al., 2009).
The ONC, NGC1333, USco, Collinder 69 and Ophiuchus all contain several massive B stars providing strong radiation and shockwaves. They are also all significantly denser than Taurus. Star formation occurs more clustered in these four regions, as opposed to Taurus where it is more distributed and occurs in isolated or small groups (Joncour et al., 2018; Galli et al., 2019). The presence of wide FFP and BD binaries in these regions therefore challenges the picture mentioned above, which suggests that higher stellar densities and/or stronger feedback from massive stars prevent or disrupt low-mass BD and FFP binaries, particularly the wide, low-binding-energy systems. Interestingly, 2MASS J0422+1934 (Todorov et al., 2014) is in the immediate vicinity of T Tau, a region that is not particularly dense but including some of the most massive members of the TMC.
However, most of these wide systems have not yet been confirmed as physical pairs and, given the stellar densities of these clusters, they could instead be chance projections of two unbound members (Akter & Goodwin, 2019) or even unrelated background sources. Confirmation of their nature is therefore necessary before drawing any firm conclusion.
@@ 6 Conclusion
We present high-spatial-resolution observations using HST and VLT of a sample of 38 FFPs and 14 low-mass BD members of the USco association, as well as 25 FFP members of the Taurus association. We detect one binary candidate in Taurus with a separation of 111.90.4 mas corresponding to 18 au at 160 pc, a mass ratio of 0.9 and estimated component masses in the range 3–6 MJup for the primary and 2.6–5.2 MJup for the secondary, according to Chabrier et al. (2023) models and assuming no extinction. No binary candidates were identified in USco and no wide binary candidates were identified in either association. Using advanced PSF-fitting methods, we also report the tentative detection of four very faint companion candidates in the F850LP filter, with flux ratios of only 3% to 5%. Given the uncertain nature of these detections, which are at the resolution and/or detection limits of the HST images, these candidates should be considered with caution and require confirmation.
@@
When combined with previous high-spatial-resolution surveys of the very low-mass populations in these regions, our findings reveal a statistically significant difference in the multiplicity properties below 25 MJup , with a binary fraction of 4.9% in Taurus and an upper limit of 1.2% in USco.
On one hand, the absence of FFP binaries in USco and their presence in Taurus suggest that dynamical interactions, higher stellar densities, and/or feedback from massive stars in USco likely prevent the formation or disrupt fragile binary systems at the lowest masses.
On the other hand, the existence of a few BD and FFP wide binary candidates reported in the literature in USco and other dense clusters including massive stars–such as the ONC, NGC 1333, Collinder 69 and Ophiuchus–appears to contradict this picture. If their binary nature is confirmed, the presence of such wide systems in these regions suggests that either some FFP binaries can survive despite strong dynamical processing (Portegies Zwart & Hochart, 2024; Huang et al., 2024) or that certain conditions—such as lower-density subregions within these clusters—allow them to persist.
@@
This paradox highlights the sensitivity of low-mass BD and FFP formation to inborn and evolutionary effects. It raises new questions about the exact role of environmental factors in shaping the low-mass BD and FFP population and underlines the need for further observational confirmation of the physical nature of these wide systems. Additionally, future observations at higher spatial resolution, particularly targeting the unexplored separation range below 7 au, will be crucial to providing a more complete picture of how environmental factors contribute to the formation and evolution of low-mass objects.
@@ Acknowledgements.
We are grateful to Tricia Royle for her support with the HST observations. JO acknowledge financial support from ”Ayudas para contratos postdoctorales de investigación UNED 2021”. DB and NH have been funded by grant No. PID2023-150468NB-I00 by the Spain Ministry of Science, Innovation/State Agency of Research MCIN/AEI/ 10.13039/501100011033 and by “ERDF A way of making Europe”. PABG acknowledges financial support from the São Paulo Research Foundation (FAPESP) under grant 2020/12518-8. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France. The original description of the VizieR service was published in Ochsenbein et al. (2000). This research made use of Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2025). This work made use of GNU Parallel (Tange, 2011), astropy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018), Topcat (Taylor, 2005), matplotlib (Hunter, 2007), Plotly (Plotly Technologies Inc., 2015), Numpy (Harris et al., 2020), aplpy (Robitaille & Bressert, 2012; Robitaille, 2019).References
- Akter & Goodwin (2019) Akter, S. & Goodwin, S. P. 2019, MNRAS, 488, 3446
@@
- De Furio et al. (2025) De Furio, M., Meyer, M. R., Greene, T., et al. 2025, ApJ, 981, L34
@@
- Miret-Roig et al. (2022) Miret-Roig, N., Bouy, H., Raymond, S. N., et al. 2022, Nature Astronomy, 6, 89
@@
- Portegies Zwart & Hochart (2024) Portegies Zwart, S. & Hochart, E. 2024, SciPost Astronomy, 3, 001
@@
- Wang et al. (2015a) Wang, J. J., Ruffio, J.-B., De Rosa, R. J., et al. 2015a, pyKLIP: PSF Subtraction for Exoplanets and Disks, Astrophysics Source Code Library, record ascl:1506.001
- Wang et al. (2015b) Wang, L., Kouwenhoven, M. B. N., Zheng, X., Church, R. P., & Davies, M. B. 2015b, MNRAS, 449, 3543
@@
| J041514.73+280009.0 | 63.81145 | 28.00243 | 22 | 19.88 | 0.05 | 18.13 | 0.05 | 18.67 | 0.01 | 17.54 | 0.01 | 15.07 | 0.05 |
| J041546.22+280811.4 | 63.94261 | 28.13649 | 6 | 26.26 | 0.21 | 23.47 | 0.05 | 24.77 | 0.21 | 23.7 | 0.25 | 19.96 | 0.14 |
| J041548.09+291132.9 | 63.95039 | 29.19245 | 23 | 20.62 | 0.05 | 18.92 | 0.05 | 19.52 | 0.01 | 18.34 | 0.01 | 15.7 | 0.05 |
| J041947.39+281534.6 | 64.94751 | 28.25954 | 7 | 23.02 | 0.05 | 21.53 | 0.05 | 22.27 | 0.04 | 21.05 | 0.04 | 18.15 | 0.05 |
| Upper Scorpius | |||||||||||||
| J154540.24-242207.2 | 236.41761 | -24.36874 | 8 | 23.41 | 0.1 | 21.99 | 0.07 | 22.9 | 0.06 | 21.9 | 0.07 | 18.55 | 0.06 |
| J154915.32-244139.1 | 237.31381 | -24.69425 | 8 | 22.92 | 0.05 | 21.27 | 0.05 | 22.07 | 0.03 | 21.04 | 0.04 | 18.2 | 0.06 |
| J155150.21-213457.4 | 237.95917 | -21.58269 | 9 | 23.09 | 0.05 | 21.28 | 0.05 | 22.27 | 0.04 | 21.23 | 0.04 | 18.21 | 0.06 |
| J155416.68-263018.1 | 238.56942 | -26.50513 | 9 | 22.27 | 0.05 | 20.62 | 0.05 | 21.5 | 0.02 | 20.46 | 0.02 | 17.58 | 0.05 |
| Object ID | RA (J2000) | Dec (J2000) | Mass | i | z | J | strehl | |||
| (deg) | (deg) | (MJup ) | (mag) | (mag) | (mag) | (mag) | (mag) | (mag) | (%) | |
| 793 | 243.58935 | -23.65409 | 21 | 18.63 | 0.05 | 17.37 | 0.05 | 14.99 | 0.05 | 2 |
| 924 | 244.96433 | -22.69256 | 28 | 18.45 | 0.05 | 17.03 | 0.05 | 14.42 | 0.05 | 3 |

### v2

@@ 5.3 Implications for the formation process of BDs and FFPs
This interpretation is further supported by the discovery of wide binary BD and FFP candidates in several young, nearby, and relatively massive star-forming regions, where warmer initial conditions are expected. In the Orion Nebula Cluster (ONC), a dense and massive environment with an age of 1 Myr, Luhman (2024) reported one candidate binary FFP ([LAB2024] 139) among a sample of 242 brown dwarfs and planetary-mass objects. This system has a wide projected separation of 132 au and a small mass ratio of . Similarly, Langeveld et al. (2024) identified two binary candidates among six FFP candidates in the 1–3 Myr NGC 1333 cluster. Both systems (NGC1333-NN10 and NGC1333-NN12) exhibit wide separations of 164 au and 76 au, respectively, with relatively high mass ratios of –0.8.
If confirmed as physical pairs, these systems would join the small population of young, wide binary systems with components at or near the planetary-mass regime, including Oph 11 ( au; Close et al., 2007), Oph 98 ( au; Fontanive et al., 2020), the extremely wide LOri 167 system in the Collinder 69 cluster ( au; Barrado Y Navascués et al., 2007), and the benchmark 2MASS J1207-3932 system in the TW Hya association (Chauvin et al., 2004), illustrated in Fig. 11. Similar weakly bound and slightly more massive wide BD binary candidates have also been reported in Upper Sco (Bouy et al., 2006; Béjar et al., 2008) and in Orion (De Furio et al., 2022). In contrast, only one wide BD binary has been reported to date in Taurus (FU Tau B; Luhman et al. 2009), which is coincidentally located in the vicinity of -Tau (B9V) and 62 Tau (B3V), two of the very few B-stars present in Taurus. However, most of these wide systems have not yet been confirmed as physical pairs and, given the stellar densities of these clusters, they could instead be chance projections of two unbound members or even unrelated background sources. Confirmation of their nature is therefore necessary before drawing any firm conclusion. But if confirmed, the existence of wide systems in massive and dynamically active regions and lack thereof in Taurus is qualitatively consistent with expectations from fragmentation in warmer parental clouds, where larger Jeans lengths naturally lead to wider initial separations. In this context, the lack of detected binaries below 30–50 MJup in Upper Sco may not reflect an absence of multiplicity altogether, but rather the absence of close and moderately bound systems in this mass regime. Low-mass binaries formed with intrinsically wide separations would be more difficult to detect and more susceptible to disruption, naturally reducing the observed binary fraction at small separations in warmer clouds.
Finally, dynamical evolution may further accentuate these differences. Even if wide, low-mass binaries form efficiently in dense or massive associations, their low binding energies make them particularly vulnerable to early disruption through interactions with other members or with the evolving cluster potential. As a result, the observed multiplicity properties in Upper Sco likely reflect a combination of initial formation conditions and subsequent dynamical processing, whereas the colder and sparser environment of Taurus may favor the formation and long-term survival of closer, bound systems in the planetary-mass regime.
@@ 6 Conclusion
We present high-spatial-resolution observations using HST and VLT of a sample of 38 FFPs and 14 low-mass BD members of the USco association, as well as 25 FFP members of the Taurus association. We detect one binary candidate in Taurus with a separation of 111.90.4 mas corresponding to 18 au at 160 pc, a mass ratio of 0.9 and estimated component masses in the range 3–6 MJup for the primary and 2.6–5.2 MJup for the secondary, according to Chabrier et al. (2023) models and assuming no extinction. Spectroscopic observations are required to confirm the nature of this pair. No binary candidates were identified in USco and no wide binary candidates were identified in either association. Using advanced PSF-fitting methods, we also report the tentative detection of four very faint companion candidates in the F850LP filter, with flux ratios of only 3% to 5%. Given the uncertain nature of these detections, which are at the resolution and/or detection limits of the HST images, these candidates should be considered with caution and require confirmation.
@@
When combined with previous high-spatial-resolution surveys of the very low-mass populations in these regions, our findings reveal a statistically significant difference in the multiplicity properties below 30–50 MJup , with a binary fraction of % below 29MJup and % below 50MJup in Taurus and an upper limit of 1% in USco over the corresponding luminosity range.
The absence of visual binaries below 30–50 MJup in Upper Sco, contrasted with several detections in Taurus over the same mass range, may reflect intrinsically different formation conditions rather than observational biases alone. Upper Sco and Taurus probe markedly distinct star-forming environments: Taurus formed in cold, sparse molecular filaments, whereas Upper Sco originated in a more massive, warmer cloud complex influenced by early massive stars. In such warmer environments, the Jeans mass is expected to increase and the characteristic fragmentation scale to shift toward larger separations, naturally suppressing the formation of the lowest-mass bound systems and favoring wider, more weakly bound pairs. This interpretation is qualitatively consistent with the existence of a few wide brown dwarf and planetary-mass binaries reported in Upper Sco, Orion, and other massive associations, and with the relative scarcity of close, low-mass systems in these regions. While dynamical disruption may further reduce the survivability of the most weakly bound binaries, especially in dense clusters, the lack of close companions at the lowest masses in Upper Sco is likely rooted primarily in the initial fragmentation process. Although speculative at this stage, this environmentally driven picture provides a physically motivated framework to interpret the observed multiplicity differences and can be tested with future spectroscopic confirmation of candidates, improved constraints on cloud temperatures, and deeper high-resolution surveys across a broader range of star-forming environments.
@@
These results highlight the sensitivity of low-mass BD and FFP formation to inborn and evolutionary effects. They raise new questions about the exact role of environmental factors in shaping the low-mass BD and FFP population and underlines the need for further observational confirmation of the physical nature of these wide systems. Additionally, future observations at higher spatial resolution, particularly targeting the unexplored separation range below 7 au, will be crucial to providing a more complete picture of how environmental factors contribute to the formation and evolution of low-mass objects.
@@ Acknowledgements.
We thank the referee for a prompt and thorough review that has significantly improved the quality and clarity of the manuscript. We are grateful to Tricia Royle for her support with the HST observations. Based in part on observations with the NASA/ESA Hubble Space Telescope obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for Program number HST-GO-17167 was provided through a grant from the STScI under NASA contract NAS5-26555. A.P. acknowledges financial support from the UNAM-PAPIIT IN120226 grant, México. JO acknowledge financial support from ”Ayudas para contratos postdoctorales de investigación UNED 2021”. DB and NH have been funded by grant No. PID2023-150468NB-I00 by the Spain Ministry of Science, Innovation/State Agency of Research MCIN/AEI/ 10.13039/501100011033 and by “ERDF A way of making Europe”. PABG acknowledges financial support from the São Paulo Research Foundation (FAPESP) under grant 2020/12518-8. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France. The original description of the VizieR service was published in Ochsenbein et al. (2000). This research made use of Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2025). This work made use of GNU Parallel (Tange, 2011), astropy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018), Topcat (Taylor, 2005), matplotlib (Hunter, 2007), Plotly (Plotly Technologies Inc., 2015), Numpy (Harris et al., 2020), aplpy (Robitaille & Bressert, 2012; Robitaille, 2019).References
- Alves et al. (2025) Alves, J., Lombardi, M., & Lada, C. J. 2025, A&A, 697, A208
@@
- De Furio et al. (2025b) De Furio, M., Meyer, M. R., Greene, T., et al. 2025b, ApJ, 981, L34
- De Furio et al. (2022) De Furio, M., Meyer, M. R., Reiter, M., et al. 2022, ApJ, 925, 112
- del Burgo & Laureijs (2005) del Burgo, C. & Laureijs, R. J. 2005, MNRAS, 360, 901
- Delorme et al. (2010) Delorme, P., Albert, L., Forveille, T., et al. 2010, A&A, 518, A39
@@
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
@@
- Luhman et al. (2023) Luhman, K. L., Tremblin, P., Birkmann, S. M., et al. 2023, The Astrophysical Journal Letters, 949, L36
- Martín et al. (2004) Martín, E. L., Delfosse, X., & Guieu, S. 2004, AJ, 127, 449
@@
- Miret-Roig et al. (2022a) Miret-Roig, N., Bouy, H., Raymond, S. N., et al. 2022a, Nature Astronomy, 6, 89
- Miret-Roig et al. (2022b) Miret-Roig, N., Galli, P. A. B., Olivares, J., et al. 2022b, A&A, 667, A163
@@
- Schmitt et al. (2022) Schmitt, J. H. M. M., Czesla, S., Freund, S., Robrade, J., & Schneider, P. C. 2022, A&A, 661, A40
@@
- Wang et al. (2015) Wang, J. J., Ruffio, J.-B., De Rosa, R. J., et al. 2015, pyKLIP: PSF Subtraction for Exoplanets and Disks, Astrophysics Source Code Library, record ascl:1506.001
@@
- Žerjal et al. (2023) Žerjal, M., Ireland, M. J., Crundall, T. D., Krumholz, M. R., & Rains, A. D. 2023, MNRAS, 519, 3992
@@
| J041514.73+280009.0 | 63.81145 | 28.00243 | 0.021 | 19.88 | 0.05 | 18.13 | 0.05 | 18.67 | 0.01 | 17.54 | 0.01 | 15.07 | 0.05 |
| J041546.22+280811.4 | 63.94261 | 28.13649 | 0.005 | 26.26 | 0.21 | 23.47 | 0.05 | 24.77 | 0.21 | 23.7 | 0.25 | 19.96 | 0.14 |
| J041548.09+291132.9 | 63.95039 | 29.19245 | 0.021 | 20.62 | 0.05 | 18.92 | 0.05 | 19.52 | 0.01 | 18.34 | 0.01 | 15.7 | 0.05 |
| J041947.39+281534.6 | 64.94751 | 28.25954 | 0.007 | 23.02 | 0.05 | 21.53 | 0.05 | 22.25 | 0.04 | 21.14 | 0.04 | 18.15 | 0.05 |
| Upper Sco | |||||||||||||
| J154540.24-242207.2 | 236.41761 | -24.36874 | 0.008/0.010 | 23.41 | 0.1 | 21.99 | 0.07 | 22.9 | 0.06 | 21.9 | 0.07 | 18.55 | 0.06 |
| J154915.32-244139.1 | 237.31381 | -24.69425 | 0.008/0.010 | 22.92 | 0.05 | 21.27 | 0.05 | 22.07 | 0.03 | 21.04 | 0.04 | 18.2 | 0.06 |
| J155150.21-213457.4 | 237.95917 | -21.58269 | 0.008/0.011 | 23.09 | 0.05 | 21.28 | 0.05 | 22.27 | 0.04 | 21.23 | 0.04 | 18.21 | 0.06 |
| J155416.68-263018.1 | 238.56942 | -26.50513 | 0.009/0.015 | 22.27 | 0.05 | 20.62 | 0.05 | 21.5 | 0.02 | 20.46 | 0.02 | 17.58 | 0.05 |
| Object ID | RA (J2000) | Dec (J2000) | i | z | J | strehl | Mass | |||
| (deg) | (deg) | (mag) | (mag) | (mag) | (mag) | (mag) | (mag) | (%) | (M⊙) | |
| 793 | 243.58935 | -23.65409 | 18.63 | 0.05 | 17.37 | 0.05 | 14.99 | 0.05 | 2 | 0.02/0.039 |
| 924 | 244.96433 | -22.69256 | 18.45 | 0.05 | 17.03 | 0.05 | 14.42 | 0.05 | 3 | 0.027/0.056 |
| 1553 | 241.46574 | -22.10589 | 17.32 | 0.05 | 16.15 | 0.05 | 13.91 | 0.05 | 31 | 0.029/0.052 |
| 768 | 242.76503 | -24.44526 | 19.06 | 0.05 | 17.51 | 0.05 | 14.72 | 0.05 | 8 | 0.025/0.03 |
@@ Appendix A HST Photometry
Appendix B Summary of literature surveys
| Name | RA (J2000) | Dec (J2000) | i | z | SpType | Ref. SpType | Multiple? | Ref. Mult. | ||
| (deg) | (deg) | (mag) | (mag) | (mag) | (mag) | |||||
| V928 Tau | 68.07858 | 24.37419 | M0.8 | (4) | Yes | (2) | ||||
| 2MASS J04414565+2301580 | 70.44022 | 23.0328 | 13.366 | 0.05 | 12.517 | 0.05 | M4.3 | (4) | Yes | (1) |
| 2MASS J04390163+2336029 | 69.7568 | 23.60083 | 13.678 | 0.001 | 12.925 | 0.006 | M4.9 | (4) | No | (1),(2) |
| 2MASS J04554757+3028077 | 73.94822 | 30.46881 | 13.925 | 0.003 | 13.001 | 0.009 | M4.75 | (4) | Yes | (1),(2) |
| 2MASS J04354526+2737130 | 68.9386 | 27.6203 | 19.757 | 0.032 | 18.043 | 0.025 | M9.25 | (1) | No | (1) |
| 2MASS J04151471+2800096 | 63.81131 | 28.00267 | 19.893 | 0.016 | 18.15 | 0.014 | M8.5 | (4) | No | (1) |
| 2MASS J04355143+2249119 | 68.9643 | 22.81999 | 20.428 | 0.043 | 18.697 | 0.021 | M8.5 | (4) | No | (1) |
| 2MASS J04272799+2612052 | 66.86665 | 26.20146 | 20.443 | 0.04 | 18.727 | 0.026 | M9.5 | (4) | No | (1) |
| 2MASS J04574903+3015195 | 74.4543 | 30.25542 | 20.798 | 0.066 | 18.97 | 0.037 | M9.25 | (4) | No | (1),(2) |
| 2MASS J04215450+2652315 | 65.4771 | 26.87542 | 20.817 | 0.05 | 18.999 | 0.023 | M8.5 | (4) | No | (1),(2) |
| 2MASS J04335245+2612548 | 68.46857 | 26.21524 | 20.822 | 0.031 | 19.134 | 0.019 | M8.25 | (4) | No | (1),(2) |
| 2MASS J04190126+2802487 | 64.75529 | 28.04686 | 21.066 | 0.079 | 19.657 | 0.034 | M9 | (4) | No | (1) |
| 2MASS J04373705+2331080 | 69.40437 | 23.51891 | 22.592 | 0.05 | 21.065 | 0.169 | L0 | (4) | No | (1) |
| Name | RA (J2000) | Dec (J2000) | i | z | SpType | Ref. SpType | Multiple? | Ref. Mult. | ||
| (deg) | (deg) | (mag) | (mag) | (mag) | (mag) | |||||
| 2MASS J15565545-2258403 | 239.23108 | -22.97788 | 12.349 | 0.05 | 12.314 | 0.05 | M0.5 | (5) | No | (3) |
| 2MASS J16074449-2036030 | 241.93539 | -20.60086 | 12.732 | 0.05 | 12.12 | 0.05 | M4 | (5) | Yes | (3) |
| 2MASS J16094644-1937361 | 242.44351 | -19.6267 | 12.832 | 0.05 | 12.378 | 0.05 | M1 | (5) | No | (3) |
| 2MASS J16081843-2232248 | 242.07682 | -22.54023 | 20.821 | 0.05 | 18.994 | 0.05 | M9.25 | (5) | No | (1),(4) |
| 2MASS J16071478-2321011 | 241.81159 | -23.35031 | 21.186 | 0.05 | 19.38 | 0.05 | M9.25 | (5) | No | (1),(4) |
| 2MASS J16391914-2534093 | 249.82978 | -25.56928 | 21.227 | 0.05 | 20.777 | 0.132 | M9.5 | (5) | No | (1),(4) |
| 2MASS J16073799-2242468 | 241.90829 | -22.71301 | 21.422 | 0.05 | 19.651 | 0.05 | M9 | (5) | No | (1),(4) |
| 2MASS J16072782-2239040 | 241.86593 | -22.65113 | 21.597 | 0.05 | 19.774 | 0.05 | M9.25 | (5) | No | (1),(4) |
| 2MASS J16122895-2159358 | 243.12063 | -21.99328 | 22.026 | 0.05 | 19.891 | 0.05 | M9.5 | (5) | No | (1),(4) |
| 2MASS J16130232-2124283 | 243.25968 | -21.40788 | 22.11 | 0.05 | 20.129 | 0.05 | M9.5 | (5) | No | (1),(4) |
| UGCS J160843.43-224516.0 | 242.181 | -22.75444 | 22.379 | 0.05 | 21.033 | 0.05 | L1 | (5) | No | (1),(4) |
| UGCS J160918.67-222923.8 | 242.32788 | -22.48992 | 22.924 | 0.05 | 21.008 | 0.05 | L1 | (5) | No | (1),(4) |